OpenStudy (anonymous):

What is the sum of a 14–term arithmetic sequence where the last term is 30 and the common difference is –5? What formula do i use?

OpenStudy (anonymous):

sum of AS = (1/2) (first term + last term)(no. of terms) last term = 30 first term= 30-(-5)(14-1) no. of term=14